How many grams of solid sodium fluoride should be added to 0.500 L of a 0.137 M hydrofluoric acid solution to prepare a buffer with a pH of 2.487 ?
Answer #1

Henderson - Hasselbalch equation- IpH = pka + log (Conjugate base] I weak and pka CHF) = 3.15 12-487= 3.15 + log LF in CHFJ m